# Boosted second moment method in random regular graphs

### v1

@@ Abstract.
Determining the asymptotic independence ratio of random regular graphs is a key challenge in the area of sparse random graphs. Due to the interpolation method, very good upper bounds are known, which are actually known to be sharp for sufficiently large degrees.
This paper provides explicit lower bounds for given degrees $d$, beating the previous best bounds for $d\geq 10$. The starting point is a second moment argument that we can boost by arguing that the obtained independent set has a certain spatial Markov property. One can then exploit this property by making local modifications to the independent set, resulting in substantial improvements.
Our approach can also be used to prove the existence of other objects in random regular graphs. To demonstrate this, we consider the problem of decomposing random regular graphs into stars.
@@ 1. Introduction
This paper is concerned with random regular graphs: let $\mathbb{G}_{N,d}$ be a uniform random graph among all $d$-regular simple graphs on the vertex set $\{1,\ldots,N\}$. A standard Azuma-type argument shows that the independence ratio of $\mathbb{G}_{N,d}$ is highly concentrated around its expectation for any fixed $d$ and $N$. A classic result [4] shows that for each degree $d\geq 3$, there exists a constant $\alpha^{\star}_{d}$ such that the independence ratio of the random $d$-regular graph $\mathbb{G}_{N,d}$ converges in probability to $\alpha^{\star}_{d}$ as $N\to\infty$.
Determining $\alpha^{\star}_{d}$, or at least providing good bounds for it, is a major challenge of the area. In [11, 10] Frieze and Łuczak used a second moment argument to give accurate asymptotics for $\alpha^{\star}_{d}$ as $d\to\infty$:
(1) $\alpha^{\star}_{d}=\frac{2}{d}\bigg(\log d-\log\log d+1-\log 2+o_{d}(1)\bigg).$
@@
Non-rigorous statistical physics methods suggested that the precise value of $\alpha^{\star}_{d}$ is given by a certain 1-RSB (1-step replica symmetry breaking) formula for $d\geq 20$ [3]. In their seminal work [8] Ding, Sun, and Sly confirmed this prediction for very large $d$. Below $20$ there is a phase transition: for $d\leq 19$ one can prove upper bounds that are stronger than the 1-RSB formula [12]. These RSB bounds are likely to be extremely close to the true value.
@@
As for lower bounds for specific degrees, however, there is still a long way to go to close the gaps. Most results in this direction attained lower bounds by analyzing certain (local) processes/algorithms. [^1] Many results actually apply to the more general setting of graphs of (essentially) large girth. A noteable tool is Wormald’s differential equation method. [^2] This method was introduced in [15]. See [14] for a concise treatment. However, when the algorithms are more sophisticated, then it tends to be very challenging to analyze them, often requiring considerable computer assistance. Due to such difficulties, some approaches treat only the smallest cases $d=3$ or $d=3,4$. Currently, the best lower bounds are due to Csóka [5] for $d=3,4$; to Duckworth–Zito [9] for $d=5,6,7$; and to Wormald [15] for $d\geq 8$, apart from $d=16$ for which degree a better bound was achieved via star decompositions in [6].
In this paper we revisit the second moment method. As we mentioned, the asymptotic formula (1) for $\alpha^{\star}_{d}$ was attained using this method [10]. In fact, that result built on an earlier work [11] regarding the independence ratio of sparse Erdős–Rényi random graphs, and the second moment method was actually used in that earlier work. Our approach applies the second moment method directly to random regular graphs. Instead of an asymptotic formula, this gives us explicit lower bounds on $\alpha^{\star}_{d}$ for specific degrees $d$. These bounds would already provide new results, but we can boost them further with the following additional step. One can show that the second moment method actually proves the existence of independent sets with a certain spatial Markov property. This allows us to precisely calculate how much improvement we may obtain by making local modifications to our independent set; see Section 1.2 for details. Table 1 shows our new bounds compared to the best bounds previously known. We also included the 1-RSB upper bounds in the table (which are conjectured to be sharp for $d\geq 20$). Our lower bounds are fairly close: for $d=500$ we have $0.0190$ (lower) versus $0.0193$ (upper).
degree best previous low.bnd our low.bnd best upp.bnd
$10$ [15]$0.2573$ $0.258371$ $0.281105$
$11$ [15]$0.2447$ $0.246775$ $0.268311$
$12$ [15]$0.2335$ $0.236447$ $0.256856$
$13$ [15]$0.2234$ $0.227167$ $0.246525$
$14$ [15]$0.2143$ $0.218767$ $0.237148$
$15$ [15]$0.2061$ $0.211113$ $0.228588$
$16$ [6]$0.2000$ $0.204101$ $0.220737$
$17$ [15]$0.1916$ $0.197645$ $0.213502$
$18$ [15]$0.1852$ $0.191674$ $0.206810$
1919 [15]$0.1793$ $0.186131$ $0.200597$
2020 [15]$0.1738$ $0.180967$ $0.194809$
$50$ [15]$0.0951$ $0.104315$ $0.109797$
$100$ [15]$0.0572$ $0.065030$ $0.067446$
$200$ [15, Thm 4]$0.0260$ $0.039010$ $0.040018$
$500$ [15, Thm 4]$0.0123$ $0.019005$ $0.019307$
$1000$ [15, Thm 4]$0.0068$ $0.010771$ $0.010890$
$2000$ [15, Thm 4]$0.0037$ $0.006014$ $0.006061$
$5000$ [15, Thm 4]$0.0017$ $0.002736$ $0.002750$
$10000$ [15, Thm 4]$0.0009$ $0.001493$ $0.001498$
Table 1. Comparing bounds on the independence ratio $\alpha^{\star}_{d}$
We point out that most papers in the literature only reported their bounds for a small number of degrees. In fact, some of these papers were actually able to carry out the required computer-assisted analysis only for small degrees. For good explicit bounds for degrees $d>20$, one has had to use a weaker bound provided by an explicit formula given in [15, Theorem 4].
@@
Although our bounds do not come in the shape of a closed formula, it is relatively simple to compute the exact values with a computer. Our short SageMath code is able to quickly compute the bound for any given degree $d\leq 500,000$. We uploaded the code to the following GitHub repository.
@@
In the remainder of the introduction, we give a brief overview of our approach.

### v2

@@ Abstract.
Determining the asymptotic independence ratio of random regular graphs is a key challenge in the area of sparse random graphs. Due to the interpolation method, we have very good upper bounds at our disposal, which are actually known to be sharp for sufficiently large degrees. However, we are still in need of good explicit lower bounds for specific degrees.
The classical approach by Frieze and Łuczak achieves a lower bound by first applying the second moment method to sparse Erdős–Rényi graphs, and then cleverly transitioning from that model to regular graphs. They obtain an asymptotic formula (as the degree tends to infinity) but no explicit lower bounds are derived for specific degrees.
In contrast, in this paper, we apply the second moment method directly to random regular graphs. This approach has a number of advantages. First, we can numerically compute good explicit lower bounds for any given degree $d$. Moreover, we can even boost this lower bound by arguing that the obtained independent set has a certain spatial Markov property. One can then exploit this property by making local modifications to the independent set, resulting in substantial improvements, and beating the previous best bounds for any $d\geq 10$. Finally, this method gives finer asymptotics as $d\to\infty$ than the original Frieze–Łuczak approach.
Moreover, these results can be useful even beyond the scope of the independence ratio due to the fact that independent sets with the Markov property may be used to construct other objects in random regular graphs. To demonstrate this, we consider the problem of decomposing random regular graphs into stars.
@@ 1. Introduction
This paper is concerned with random regular graphs: let $\mathbb{G}_{N,d}$ be a uniform random graph among all $d$-regular simple graphs on the vertex set $\{1,\ldots,N\}$. A standard Azuma-type argument shows that the independence ratio [^1] The independence ratio is the size of the largest independent set divided by the total number of vertices. of $\mathbb{G}_{N,d}$ is highly concentrated around its expectation for any fixed $d$ and $N$. A classic result [4] shows that for each degree $d\geq 3$, there exists a constant $\alpha^{\star}_{d}$ such that the independence ratio of the random $d$-regular graph $\mathbb{G}_{N,d}$ converges in probability to $\alpha^{\star}_{d}$ as $N\to\infty$. Determining $\alpha^{\star}_{d}$, or at least providing good bounds for it, is a major challenge of the area.
In [5] Bollobás used random regular graphs to prove the existence of large-girth regular graphs with small independence number. His argument actually shows that $\alpha^{\star}_{d}\leq\alpha^{\mathrm{FM}}_{d}$, where the first moment bound $\alpha^{\mathrm{FM}}_{d}$ is the unique root of a certain function $\varphi_{d}\colon(0,1/2)\to\mathbb{R}$; see (7). Asymptotically, as $d\to\infty$, it is easy to show that $\alpha^{\mathrm{FM}}_{d}$ decays as
(1) $\frac{2}{d}\bigg(\log d-\log\log d+1-\log 2+o_{d}(1)\bigg).$
In the Appendix we provide sharper estimates for $\alpha^{\mathrm{FM}}_{d}$, see Theorem 11.1.
In [11, 12] Frieze and Łuczak used a second moment argument to obtain an asymptotically matching lower bound to prove that $\alpha^{\star}_{d}$ is actually asymptotically equal to (1).
@@
So we have a good understanding of the asymptotic behavior of $\alpha^{\star}_{d}$, but is it possible to determine the precise value of $\alpha^{\star}_{d}$ for a given $d$? Non-rigorous statistical physics methods [3] suggested that, for $d\geq 20$, the precise value of $\alpha^{\star}_{d}$ is given by a certain 1-RSB (1-step replica symmetry breaking) formula. In their seminal work [9] Ding, Sun, and Sly confirmed this prediction for very large dd. [^2] There is a phase transition below degree $20$: for $d\leq 19$ one can prove upper bounds that are stronger than the 1-RSB formula [14]. These RSB bounds are likely to be extremely close to the true value. Although there is no closed-form expression for this 1-RSB value, [9, Theorem 3.1] proves the following asymptotic comparison to the first moment bound:
(2) $\alpha^{\star}_{d}=\alpha^{\mathrm{FM}}_{d}-\left(\frac{2}{e}\,\frac{\log d}{d}\right)^{2}\left(1+\mathcal{O}\left(\frac{\log\log d}{\log d}\right)\right).$
@@
Note, however, that neither the Frieze–Łuczak, nor the Ding–Sly–Sun result is helpful when one is interested in $\alpha^{\star}_{d}$ for some specific degree, say, for $d=100$. What concrete lower bounds do we have at our disposal? Most results in this direction attained lower bounds by analyzing certain (local) processes/algorithms. [^3] Many results actually apply to the more general setting of graphs of (essentially) large girth. A notable tool is Wormald’s differential equation method. [^4] This method was introduced in [21]. See [19] for a concise treatment. The difficulty with this method is that it tends to be very challenging to carry out the required computer analysis for more sophisticated algorithms. Due to such difficulties, some approaches treat only the smallest cases $d=3$ or $d=3,4$. Currently, the best lower bounds are due to Csóka [6] for $d=3,4$; to Duckworth–Zito [10] for $d=5,6,7$; and to Wormald [21] for $d\geq 8$, apart from $d=16$ for which degree a better bound was achieved via star decompositions in [7].
1.1. Our bounds
In this paper we revisit the second moment method, which was used in [12] to attain the asymptotic formula (1) for $\alpha^{\star}_{d}$. In fact, [12] built on an earlier work [11], which addressed the same problem for sparse Erdős–Rényi random graphs, and the second moment method was actually used in that earlier work. Our approach applies the second moment method directly to random regular graphs. The statement of the result involves a function $f_{d,\alpha}\colon[0,\alpha]\to\mathbb{R}$ with a rather complicated but explicit formula; see (8).
Theorem 1.1.
Let $d\geq 3$ and $\alpha\in(0,1/2)$. Assume that the function $f_{d,\alpha}(\beta)$, defined in (8), has the following property:
(3) $f_{d,\alpha}\text{ attains its maximum at }\alpha^{2}.$
Then the independence ratio of $\mathbb{G}_{N,d}$ is at least $\alpha-o_{N}(1)$ with probability $1-o_{N}(1)$ as $N$ goes to $\infty$. In particular, $\alpha^{\star}_{d}\geq\alpha$.
For each degree we want to find the largest $\alpha$ for which this holds.
Definition 1.2.
For an integer $d\geq 3$ let
[…]
We refer to $\alpha^{\mathrm{SM}}_{d}$ as the second moment (lower) bound.
By Theorem 1.1 we have $\alpha^{\star}_{d}\geq\alpha^{\mathrm{SM}}_{d}$. How sharp is this bound? We will give a very precise asymptotic estimate for $\alpha^{\mathrm{SM}}_{d}$ but first let us focus on the obtained bounds for specific degrees.
We will see that verifying condition (3) essentially requires comparing two local maxima of $f_{d,\alpha}$. Therefore, it is straightforward to determine $\alpha^{\mathrm{SM}}_{d}$ with high precision using a computer for a specific degree $d$. It turns out that the lower bound $\alpha^{\mathrm{SM}}_{d}$ would already provide new results for $d\geq 20$, but we can actually boost it further with the following additional step. We will see that the second moment method proves the existence of independent sets with a certain spatial Markov property. This means that if we make “local improvements” to our independent set, then we can precisely calculate the increased density. In Section 5 we will see a simple procedure that gives the following augmented bound:
(4) $\alpha^{\mathrm{aug}}_{d}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\alpha^{\mathrm{SM}}_{d}+(1-\alpha^{\mathrm{SM}}_{d})p^{d}\frac{1+\big(1-p^{d-1}\big)^{d}}{2}\text{, where }p=\frac{1-2\alpha^{\mathrm{SM}}_{d}}{1-\alpha^{\mathrm{SM}}_{d}}.$
The density increase ($\alpha^{\mathrm{aug}}_{d}$ vs $\alpha^{\mathrm{SM}}_{d}$) is substantial for degrees up to $d\approx 50$ and noticeable up to $d\approx 1000$. For larger degrees the gain is not significant: as we will see in Section 5.1, asymptotically we have $\alpha^{\mathrm{aug}}_{d}-\alpha^{\mathrm{SM}}_{d}=d^{-2+o(1)}$, which does not get us considerably closer to (the conjectured value of) $\alpha^{\star}_{d}$.
Table 1 shows our bounds $\alpha^{\mathrm{SM}}_{d}$ and $\alpha^{\mathrm{aug}}_{d}$ compared to the best bounds previously known. We also included the RSB upper bounds in the table (which are conjectured to be sharp for $d\geq 20$). Our lower bounds are fairly close: for $d=500$ we have $0.0190$ (lower) versus $0.0193$ (upper), meaning that the “optimality rate” of our lower bound is $98.4\%$.
degree best previous low.bnd
our lower bound
$\alpha^{\mathrm{aug}}_{d}$ ($\alpha^{\mathrm{SM}}_{d}$)
best upp.bnd opt.rate
$10$ [21]$0.2573$ $0.258371$  ($0.245521$) $0.281105$ $91.91\%$
$11$ [21]$0.2447$ $0.246775$  ($0.234916$) $0.268311$ $91.97\%$
$12$ [21]$0.2335$ $0.236447$  ($0.225479$) $0.256856$ $92.05\%$
$13$ [21]$0.2234$ $0.227167$  ($0.216997$) $0.246525$ $92.14\%$
$14$ [21]$0.2143$ $0.218767$  ($0.209312$) $0.237148$ $92.24\%$
$15$ [21]$0.2061$ $0.211113$  ($0.202301$) $0.228588$ $92.35\%$
$16$ [7]$0.2000$ $0.204101$  ($0.195866$) $0.220737$ $92.46\%$
$17$ [21]$0.1916$ $0.197645$  ($0.189931$) $0.213502$ $92.57\%$
$18$ [21]$0.1852$ $0.191674$  ($0.184431$) $0.206810$ $92.68\%$
1919 [21]$0.1793$ $0.186131$  ($0.179314$) $0.200597$ $92.78\%$
2020 [21]$0.1738$ $0.180967$  ($0.174538$) $0.194809$ $92.89\%$
$50$ [21]$0.0951$ $0.104315$  ($0.102334$) $0.109797$ $95.01\%$
$100$ [21]$0.0572$ $0.065030$  ($0.064305$) $0.067446$ $96.42\%$
$200$ [21, Thm 4]$0.0260$ $0.039010$  ($0.038762$) $0.040018$ $97.48\%$
$500$ [21, Thm 4]$0.0123$ $0.019005$  ($0.018948$) $0.019307$ $98.44\%$
$1000$ [21, Thm 4]$0.0068$ $0.010771$  ($0.010753$) $0.010890$ $98.92\%$
$2000$ [21, Thm 4]$0.0037$ $0.006014$  ($0.006009$) $0.006061$ $99.25\%$
$5000$ [21, Thm 4]$0.0017$ $0.002736$  ($0.002735$) $0.002750$ $99.54\%$
$10000$ [21, Thm 4]$0.0009$ $0.001493$  ($0.001492$) $0.001498$ $99.68\%$
Table 1. Bounds on the independence ratio $\alpha^{\star}_{d}$. The last column (optimality rate) shows the relative size of our best lower bound ($\alpha^{\mathrm{aug}}_{d}$) compared to the upper bound.
@@ 1.1. Our bounds
As we mentioned, although our bounds do not come in the shape of a closed formula, it is relatively simple to obtain precise numerical values with a computer. Our SageMath code is able to quickly compute the bound for any given degree $d\leq 10^{9}$. We uploaded the code to the following GitHub repository.
@@
We point out that most papers in the literature only reported their bounds for a small number of degrees. In fact, some of them were able to carry out the required computer-assisted analysis only for the smallest degrees. Prior to this paper, in order to get an explicit lower bound on $\alpha^{\star}_{d}$ for a degree $d>20$, one has had to use a weaker bound provided by an explicit formula given in [21, Theorem 4].
To complement these numerical results, we will prove the following concrete estimate for the second moment (lower) bound $\alpha^{\mathrm{SM}}_{d}$. In the spirit of (2), we compare it to the first moment (upper) bound $\alpha^{\mathrm{FM}}_{d}$. (In turn, the reader may find accurate estimates for $\alpha^{\mathrm{FM}}_{d}$ in Section 11.)
Theorem 1.3.
Let
(5) $\varepsilon_{d}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{4\sqrt{2}}{e}\frac{(\log d)^{1/2}}{d^{3/2}}.$
Then, for any degree $d\geq 3$, it holds that
(6) $\alpha^{\mathrm{FM}}_{d}-\varepsilon_{d}\leq\alpha^{\mathrm{SM}}_{d}\leq\alpha^{\star}_{d}\leq\alpha^{\mathrm{FM}}_{d}.$
The proof will require very delicate estimates [^5] The bound $\alpha^{\mathrm{SM}}_{d}\geq\alpha^{\mathrm{FM}}_{d}-\varepsilon_{d}$ is very accurate: the inequality would fail to hold (for large $d$) if we replaced the multiplicative constant in $\varepsilon_{d}$ with anything smaller., occupying a significant portion of the paper (Sections 710).
Note that the Ding–Sly–Sun result [9] determines $\alpha^{\star}_{d}$ precisely for very large $d$, showing that $\alpha^{\star}_{d}=\alpha^{\mathrm{FM}}_{d}-1/d^{2-o(1)}$; see (2). Our lower bound (6) has the advantage that it produces an explicit bound for every $d$, although with the somewhat weaker exponent 3/2o(1)3/2-o(1).
1.2. Applications
The significance of the existence of Markovian independent sets of density $\alpha=\alpha^{\mathrm{SM}}_{d}$ goes beyond obtaining the lower bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{aug}}_{d}$. One can start from a Markovian independent set and try to create other objects, and hence proving the existence of more complex structures in random regular graphs. (For instance, suppose that we want to find large induced subgraphs with degree bound, say, $2$. Then it is not hard to design a local algorithm that produces such a subgraph from an independent set.)
In this paper we present one such application, concerning so-called star decompositions. In Section 6 we demonstrate how our approach can be used to prove that the edge set of a random $d$-regular graph asymptotically almost surely (a.a.s) [^6] For a fixed $d$, we say that $\mathbb{G}_{N,d}$ a.a.s. has a property if it holds with probability $1-o_{N}(1)$. can be decomposed into $k$-stars for certain pairs $d,k$.
We also present a consequence of our numerical bounds on $\alpha^{\star}_{d}$. There used to be a conjecture in sparse graph limit theory stating that so-called typical processes are the same as the weak limits of factor of IID processes. Loosely speaking, this says that “most optimization problems over typical, sparse graphs can be solved by local algorithms” [18]. The conjecture was refuted by a result of Gamarnik and Sudan [13] showing that a typical independent set has larger density than a factor of IID independent set for sufficiently large $d\geq d_{0}$ (with no explicit $d_{0}$ derived). Our concrete bounds show that this statement already holds true starting at degree $d=403$. The details and precise statements can be found in the Appendix, see Section 12.
Organization of the paper
  • In Section 2 we provide a proof sketch for Theorem 1.1. Later we will rigorously prove a stronger version (Theorem 4.3) building on a result of Backhausz–Bordenave–Szegedy [1] about typical processes. This section is not required for the rest of the paper, but it may help the reader to gain a clearer understanding.
  • The next part of the paper contains our “qualitative results”. Section 3 gathers the necessary facts about invariant and typical processes on the infinite tree. In Section 4 we examine the second moment condition for independent sets, while in Section 5 we discuss how the Markovian independent set can be augmented.
  • Section 6 is concerned with the application for star decompositions.
  • The remainder of the paper (the “quantitative part”) is dedicated to the proof of Theorem 1.3. Section 7 contains some general estimates. We analyze the function $f_{d,\alpha}$ and its local optima in Section 8, while Section 9 focuses on the second local maximum and contains the main steps of the proof of Theorem 1.3, with two technical arguments postponed until Section 10.
  • Finally, we included an Appendix containing sharp estimates for the first moment bound $\alpha^{\mathrm{FM}}_{d}$ (Section 11) and results about factor of IID independent sets (Section 12).
2. Counting independent sets in random regular graphs
This section is somewhat independent of the rest of the paper, as its results are not needed in what follows, and stronger statements will be established later.
2.1. Proof sketch of Theorem 1.1
Our aim here is to present a short, self-contained proof of our basic bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{SM}}_{d}$ (Theorem 1.1). The argument relies only on standard results from the literature and makes direct use of the second moment method.
In a later section, we prove a stronger result (Theorem 4.3), where the second moment method is replaced by a theorem of Backhausz–Bordenave–Szegedy concerning typical processes with a spatial Markov property. That stronger result will imply the augmented bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{aug}}_{d}$.
Recall that $\mathbb{G}_{N,d}$ denotes the uniform random graph among all $d$-regular simple graphs on the vertex set $\{1,\ldots,N\}$. This is closely related to the random multigraph $\mathbb{G}^{\textrm{conf}}_{N,d}$, produced by the so-called configuration model. Given $N$ vertices, each with $d$ “half-edges”, the configuration model picks a random pairing of these $Nd$ half-edges, resulting in $Nd/2$ edges. The corresponding random graph $\mathbb{G}^{\textrm{conf}}_{N,d}$, which is always $d$-regular, may have loops and multiple edges. A well-known fact is that if $\mathbb{G}^{\textrm{conf}}_{N,d}$ is conditioned to be simple, then we get back $\mathbb{G}_{N,d}$. Moreover, for any fixed $d$, the probability that $\mathbb{G}^{\textrm{conf}}_{N,d}$ is simple converges to a positive constant $p_{d}$ as $N\to\infty$. It follows that if $\mathbb{G}^{\textrm{conf}}_{N,d}$ a.a.s. has a certain property, then so does $\mathbb{G}_{N,d}$.
The advantage of working with $\mathbb{G}^{\textrm{conf}}_{N,d}$ is that it makes counting fairly easy: for various graph structures or objects, one can write up their expected number in $\mathbb{G}^{\textrm{conf}}_{N,d}$ using factorials (or multinomial coefficients). These quantities often have exponential growth (in $N$), the rate of which can be expressed using entropy.
In our setting, for fixed $d$ and $\alpha\in(0,1/2)$, let us define the random variable $Z_{N}$ as the number of independent sets of density $\alpha$ in $\mathbb{G}^{\textrm{conf}}_{N,d}$, and consider its first and second moments: $\mathbb{E}Z_{N}$ and $\mathbb{E}Z_{N}^{2}$.
It is fairly easy to derive the exponential rate of $\mathbb{E}Z_{N}$:
(7) $\varphi_{d}(\alpha)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=h(\alpha)+\frac{d}{2}h(1-2\alpha)-(d-1)h(1-\alpha),$
where
$h(x)=\begin{cases}-x\log x&\text{if }x\in(0,1];\\ 0&\text{if }x=0.\end{cases}$
More precisely,
$\mathbb{E}Z_{N}=(Nd)^{\mathcal{O}(1)}\exp\big(N\varphi_{d}(\alpha)\big).$
The point is that the polynomial term $(Nd)^{\mathcal{O}(1)}$ is negligible in the sense that $\varphi_{d}(\alpha)<0$ implies that $\mathbb{E}Z_{N}\to 0$ as $N\to\infty$. Since the random variable is nonnegative and integer-valued, it follows that $\mathbb{P}(Z_{N}>0)\leq\mathbb{E}Z_{N}$ also converges to 0, that is, $\mathbb{G}^{\textrm{conf}}_{N,d}$ (and hence $\mathbb{G}_{N,d}$) a.a.s. has no independent set of density $\alpha$. Recall that $\alpha^{\mathrm{FM}}_{d}$ is defined as the (unique) root of $\varphi_{d}(\alpha)$. Since $\varphi_{d}(\alpha)<0$ for every $\alpha>\alpha^{\mathrm{FM}}_{d}$, we can conclude that $\alpha^{\star}_{d}\leq\alpha^{\mathrm{FM}}_{d}$. This is the first moment (upper) bound.
As for the second moment $\mathbb{E}Z_{N}^{2}$, we need to consider pairs of independent sets (both with density $\alpha$). We partition these pairs based on the intersection size. For $\beta\in[0,\alpha]$ let $Z^{(2)}_{N,\beta}$ be the number of pairs $(A_{1},A_{2})$ where both $A_{1}$ and $A_{2}$ are independent sets of $\mathbb{G}^{\textrm{conf}}_{N,d}$ with density $\alpha$ such that their intersection $A_{1}\cap A_{2}$ has density $\beta$. It can be shown [^7] The formulas for $\varphi_{d}$ and $f_{d,\alpha}$ stem from the entropy of certain discrete distributions. Section 4 explains in detail which distributions need to be considered. that
$\mathbb{E}Z^{(2)}_{N,\beta}=(Nd)^{\mathcal{O}(1)}\exp\big(Nf_{d,\alpha}(\beta)\big),$
where, for $\beta\in[0,\alpha]$, let
(8) $\begin{split}\gamma&\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=-(1/2-\alpha)+\sqrt{(1/2-\alpha)^{2}+(\alpha-\beta)^{2}}\text{, and}\\ f_{d,\alpha}(\beta)&\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\frac{d}{2}\bigg(2h(\beta)+2h(\gamma)+4h(\alpha-\beta-\gamma)+h(1-4\alpha+2\beta+2\gamma)\bigg)\\ &\quad-(d-1)\bigg(h(\beta)+2h(\alpha-\beta)+h(1-2\alpha+\beta)\bigg).\end{split}$
Then the exponential rate of the number of pairs (with any intersection density $\beta$) is
$\varphi_{d}^{(2)}(\alpha)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\max_{\beta\in[0,\alpha]}f_{d,\alpha}(\beta).$
By the Paley–Zygmund inequality we get
$\mathbb{P}(Z_{N}>0)\geq\frac{\big(\mathbb{E}Z_{N}\big)^{2}}{\mathbb{E}Z_{N}^{2}}=(Nd)^{\mathcal{O}(1)}\exp\big(N\big(2\varphi_{d}(\alpha)-\varphi_{d}^{(2)}(\alpha)\big)\big).$
Note that we always have
$\varphi_{d}^{(2)}(\alpha)\geq f_{d,\alpha}(\alpha^{2})=2\varphi_{d}(\alpha).$
If $f_{d,\alpha}$ attains its maximum at $\alpha^{2}$, then we have equality, and obtain that
(9) $\mathbb{P}(Z_{N}>0)>\frac{1}{(Nd)^{K}}\text{ for some constant }K.$
In other words, the independence ratio of $\mathbb{G}^{\textrm{conf}}_{N,d}$ is at least $\alpha$ with probability $\geq(Nd)^{-K}$. Our goal, however, is to show that this holds a.a.s., that is, with probability $1-o_{N}(1)$, as $N\to\infty$. Next we show how $(Nd)^{-K}$ can be turned into $1-o_{N}(1)$ at the expense of slightly lowering $\alpha$.
As we mentioned in the introduction, the independence ratio of $\mathbb{G}_{N,d}$ is tightly concentrated around its expectation. We will use the following precise statement that readily follows from [20, Theorem 2.19], which is a general Azuma-type concentration result. Let $\alpha(G)$ denote the independence ratio of $G$; then
(10) $\mathbb{P}\bigg(\big|\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})-\mathbb{E}\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\big|\geq t\bigg)\leq 2\exp\left(\frac{-Nt^{2}}{d}\right).$
We set $t=t_{N}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\log N)/\sqrt{N}$ so that
$\exp\left(\frac{-Nt^{2}}{d}\right)=\exp\left(\frac{-(\log N)^{2}}{d}\right)=o\left(\frac{1}{(Nd)^{K}}\right)\text{ for any fixed $d$ and $K>0$.}$
Since the event $Z_{N}>0$ is equivalent to $\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\alpha$, from (9) and (10) we obtain
$\mathbb{E}\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\alpha-t_{N}.$
Using (10) again, we can conclude that with probability $1-o_{N}(1)$ we have
$\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\mathbb{E}\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})-t_{N}\geq\alpha-2t_{N}.$
Since $t_{N}=o_{N}(1)$, the statement of Theorem 1.1 follows.